OpenStudy (anonymous):

which is the equation of the same line as y=3x-8 3x-2y=8 -3x-2y=-8 6x-y=16 6x-2y=16

OpenStudy (anonymous):

try moving the equation around. for example. y-3x = -8 and 3x = -8-y remember you can do 2(y=3x-8) because that is multiplying by 2 on both sides :)

OpenStudy (mertsj):

The last one